OpenStudy (anonymous):

what is the inverse of the function f(x)=x/(x+3)

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

in terms of f(x)

hero (hero):

Let y = f(x) \[y = \frac{x}{x+3}\] Then swap x and y: \[x = \frac{y}{y+3}\] Then isolate y

OpenStudy (anonymous):

change f(x) by x and x by f(x)

hero (hero):

Finally after isolating y, replace it with \[f^{-1}(x)\]

OpenStudy (anonymous):

I tried that and ended up with \[y=xy+3x\] and i dont know where to go from there

OpenStudy (anonymous):



hero (hero):

\(y - xy = 3x\) \(y(1-x) = 3x\) \[y = \frac{3x}{1-x}\]

OpenStudy (anonymous):

Ah thanks! I get it, you have to pull out a factor

hero (hero):

\[f^{-1}(x) = \frac{3x}{1-x}\]